OpenStudy (anonymous):

Bag A contains 3 white marbles and 2 red marbles. Bag B contains 6 white marbles and 3 red marbles. A person draws one marble from each bag. Find the probability that both marbles are white.

OpenStudy (zarkon):

what is the probability the marble from bag A is white?

OpenStudy (anonymous):

3/5

OpenStudy (zarkon):

what about bag B...what is the prob of a white?

OpenStudy (anonymous):

6/9=2/3

OpenStudy (zarkon):

now multiply those two numbers

OpenStudy (zarkon):

assuming the draws are independent.

OpenStudy (anonymous):

2/3 * 3/5 = 6/15 = 2/5

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

Oh okay! Well thank you:)

OpenStudy (zarkon):

just so you see... P(both marbles are white)=P(white from A and white from B) =P(white from A)*P(white from B)=\(\displaystyle\frac{3}{5}\times\frac{6}{9}=\frac{2}{5}\)

OpenStudy (anonymous):

Ohh, thanks again. Looking back I don't see why I didn't get this..